[Koji EricQ]
The both arms have been locked with IR and aligned by ASS.
The IFO was left with ITMX/Y, ETMX/Y, BS, and PRM aligned, and the PSL shutter closed.
YARM
SIGNAL PATH: POY11I(+45dB)->YARM(G=+1.0)->ETMY
NORM: TRYx10
TRIG: TRY 0.01up/0.005down
FM TRIG: FM2/3/6/7/8/9 0.01up/0.05down, 0.5 sec delay
XARM
SIGNAL PATH: POX11I(+45dB)->XARM(G=+4.0)->ETMX
NORM: TRXx10
TRIG: TRX 0.01up/0.005down
FM TRIG: FM2/3/6/7/8/9 0.01up/0.05down, 0.5 sec delay
For decent locks, it was necessary that the offset of the error signals are trimmed at the input filters
even after running LSCoffset.py script.
Once the cavities were aligned for the IR, we could see the green beams are also flashing.
The Y arm was actually locked with the green with a TEM00 mode |